OpenStudy (anonymous):

Please help sketch the graph of the continuous function f:R\{-1} -> R : x-> (x+1)(X-2)/(x+1). justify your answer.

OpenStudy (anonymous):

I don't understand what that says. y = (x+1)(x-2)/(x+1) is discontinuous at x = -1 because of division by 0. You can simplify that to: \[y = (x-2) , \ x\neq -1\] Just graph that line, and put an open circle at x = -1

OpenStudy (anonymous):

when i differentiate the function i keep getting -1 as my root which i don't think is possible as you cannot divide by zero on the bottom. i'm really confused

OpenStudy (anonymous):

\[y=(x-2)\]\[dy/dx=\frac{ (x+1)(x-2)-(x+1)(2x-1) }{ (x+1)^2}\] \[\frac{ x+2 }{ (x+1) }=\frac{ 2x-1 }{ (x+1) }\] \[x+1=2x-1\] \[x=1\] but this is not necessary no maxima required

OpenStudy (anonymous):

I think you made a little error: 

OpenStudy (anonymous):

sorry

OpenStudy (phi):

when i differentiate the function i keep getting -1 as my root The roots of this "polynomial" are where it crosses the x-axis i.e. where y=0 (x+1)(X-2)/(x+1) =0 x=- 1 must be excluded because that would cause division by 0. x=2 is a valid root. If you differentiate and set that to zero you can find the min or max of the polynomial in this case you find the differential is 1, a constant. That means no max or min. Of course, if you first simplify by cancelling (x+1), you get y= x-2 which is the equation of a line with slope 1 and y-intercept -2 that would be the plot, but with an excluded value at x= -1 to account for the division by (x+1) in the original expression.